# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ Abstract.
AI systems have been known to amplify biases in real world data. Explanations may help human-AI teams address these biases for fairer decision-making. Typically, explanations focus on salient input features. If a model is biased against some protected group, explanations may include features that demonstrate this bias, but when biases are realized through proxy features, the relationship between this proxy feature and the protected one may be less clear to a human. In this work, we study the effect of the presence of protected and proxy features on participants’ perception of model fairness and their ability to improve demographic parity over an AI alone. Further, we examine how different treatments—explanations, model bias disclosure and proxy correlation disclosure—affect fairness perception and parity. We find that explanations help people detect direct biases but not indirect biases. Additionally, regardless of bias type, explanations tend to increase agreement with model biases. Disclosures can help mitigate this effect for indirect biases, improving both unfairness recognition and the decision-making fairness. We hope that our findings can help guide further research into advancing explanations in support of fair human-AI decision-making.
@@
ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learning
@@ 1. Introduction
Improving the fairness and trustworthiness of AI systems is often cited as a goal of explainable AI (XAI) (e.g., Doshi-Velez and Kim, 2017; Lipton, 2018; Barredo Arrieta et al., 2020; Das and Rad, 2020; Langer et al., 2021; Du et al., 2021; Warner and Sloan, 2021). Research in XAI aims to improve fairness in human-AI decision-making by providing insights into model predictions, and thereby allowing humans to understand and correct for model biases. On the other hand, in the context of human-AI decision-making, previous work has noted that humans often over-rely on AI predictions, and explanations can exacerbate this concern (Buçinca et al., 2021). This is especially troubling if the underlying model contains systematic biases, which may go unnoticed even when teamed with a human. In order for the human-AI team to be successful, the human needs to be able to determine when to rely on or override potentially biased AI predictions. Previous work has shown that explanations can help human-AI teams alleviate model biases when those biases depend directly on protected attributes (Dodge et al., 2019; Schoeffer et al., 2022), but little is known in the very common case that protected attributes are not explicitly included, and rather the features used for prediction contain proxies thereof (e.g., zip code for race, length of credit for age, and university attended for gender). In particular, it may be difficult for humans to identify and resolve biased model predictions based on the proxy features present in real-world data, even when explanations are provided.
In this work, we study whether explanations can help people to identify model biases and to calibrate their reliance on an AI model based on these biases. We extend this line of investigation beyond direct biases that are revealed through the use of protected (i.e., sensitive) features by considering the effect of explanations when indirect bias is revealed through proxy features which may be less obvious to a human. Further, we examine whether explicitly disclosing model biases and correlations between the proxy and protected features can help humans calibrate their trust in a biased model. Our study aims to evaluate whether explanations can directly help notice model biases, even when the biases are obfuscated by the presence of proxy features and whether explanations can help users correct model biases when they are known to be present, through the use of bias disclosure and correlation disclosure. We study the effect of these treatments (explanations, model bias disclosure and proxy correlation disclosure) on the fairness, including fairness perception and fairness in decision-making (measured by group-wise parity), as well as the accuracy of the decisions made by human-AI teams.
@@ 2. Background and Related Work
Both models and humans can be biased. Humans have been known to exhibit many implicit and unconscious biases (Greenwald and Banaji, 1995). For instance,  Bertrand and Mullainathan (2004) find that an applicant with a “White-sounding name” on a resume that is otherwise identical to a resume with an “African-American-sounding name” is more likely to receive an interview callback.
@@
As self-reported perceptions do not always align with observed behaviors in human-AI decision-making (Buçinca et al., 2020; Papenmeier et al., 2022), recent work has begun to expand out of fairness perceptions and into observed fairness in decision-making (Schoeffer et al., 2022; Wang et al., 2023). Schoeffer et al. (2022) study how explanations can help users appropriately rely on potentially unfair AI predictions.They find that explanations that highlight protected features negatively affect fairness perceptions and that decreases in fairness perception are associated with an increase in overrides of AI predictions, even on examples where this override is detrimental to the fairness of the human-AI team. Wang et al. (2023) consider the effects of the level of model bias and the presence of explanation on the fairness of human decisions. They find that explanations lead participants to make more unfair decisions, even when participants were no longer given access to model predictions or explanations.
@@ 3. Research Questions
We study the effect of explanations in improving the fairness of decisions made by human-AI teams when bias stems from different kinds of features and when participants are given different kinds of information about the model and its training data. In our study, model biases can be direct: stemming from the protected feature (gender), or indirect: stemming from a proxy feature (university) that is correlated with the protected feature. Participants may receive disclosures about the level of model bias and the strength of correlation between the proxy and protected feature. Our study addresses the following research questions:
@@
RQ5: How does dispositional trust (“an individual’s enduring tendency to trust automation” (Hoff and Bashir, 2015; Marsh and Dibben, 2003)) affect decision-making and fairness perceptions when working with models exhibiting direct or indirect bias?
RQ6: How do explanations and disclosures affect self-reported learned trust (based on “past experience or the current interaction” (Hoff and Bashir, 2015; Marsh and Dibben, 2003)) in models exhibiting direct or indirect bias?
@@ 4. Study Design
To answer the research questions posed in §3, we study decisions made by human-AI teams. [^2] This study design is IRB approved. (#$1941548$-$2$), In this study, the AI teammate is a classification model trained on partially-synthetic data in the context of loan prediction. We choose the task of loan prediction from a micro-lending platform as it is a decision-making task performed by laypeople which means that crowd-workers are more likely to have intuitions about the task and the features used in predictions. In our study, participants are shown either the protected feature of binary gender [^3] We only consider binary gender in this study. Since each participant sees only a handful of examples per task phase, it would be difficult to both show the participants a statistically realistic number of non-binary applicants and get a good sense of how participants handle anti-trans model bias. We leave this for future work. or the proxy feature of university.
{subcaptiongroup} Refer to captionRefer to caption For the applicant profile (above), an AI system has predicted that the applicant will complete the loan on time. The figure shows the weights assigned to different attributes of the applicant’s profile by the AI system.
Do you think that the applicant will be Late or will Complete the loan on time? $\circ$ Definitely Late  $\circ$ Probably Late  $\circ$ Neutral $\circ$ Probably Complete  $\circ$ Definitely Complete
@@
4.1. Conditions
In our study, we vary conditions based on the directness of bias, whether explanations are shown, and the kind of disclosure the participant receives.
  • Protected or Proxy: Whether the participant is shown a model and features including the protected (gender) or proxy (university) feature.
  • Explanations: Whether the participant is shown an input-influence explanation of how features contributed to the AI’s prediction (Figure 1, top).
  • Disclosure: In the case of direct bias, we show the participant a bias disclosure, which reveals the demographic parity (described in §5.1) of the system. In the case of indirect bias, the participant may be shown only the bias disclosure or a full bias and correlation disclosure which additionally explains the relationship between university and gender features (Figure 3).
In this study, we consider six conditions. In the first three conditions, we do not show participants explanations. Here, we consider one Protected condition with bias disclosure, and two Proxy conditions: one Proxy with correlation disclosure and one Proxy without correlation disclosure. We similarly consider three conditions with explanations allocating the biased feature and disclosure-types in the same fashion.
We assess the effect of explanations (RQ1a) comparing conditions with and without explanations (before any disclosures) in a between-subjects analysis. Similarly, we assess the effect of disclosures without explanations (RQ2a) and the effect of disclosures with explanations (RQ3a) in a within-subject analysis comparing fairness perceptions and human-AI decision-making pre- and post-disclosures. This allows us to study how disclosures may help participants in identifying model biases over what is apparent before any disclosures. These first three effects are summarized in §5.1.3. Lastly, we assess the effect of explanations and disclosures jointly (RQ4a) by comparing conditions in which participants are not shown explanations pre-disclosures with conditions in which participants are shown explanations post-disclosures. These experiments are repeated both for protected and proxy conditions to assess the differences in interventions therein (RQ1-4b).
4.2. Procedure
Our study procedure consists of three surveys (S0, S1, S2), one tutorial and warm-up phase (P0), two task phases (P1, P2), and a disclosure interlude (D) ordered as shown in §4.2.
[…]
Task Phases
In each task phase (P1, P2), the participant is shown 10 profiles of loan applicants: their features and the overall AI prediction. Depending on the condition, they may or may not be shown an explanation of the AI prediction (Figure 1 left and right, respectively). This profile will, according to the condition, include either a “gender” or a “university” feature but not both. Participants are asked to mark on a five-point-scale whether they think the applicant will complete their loan on time or be late in repaying their loan (Figure 1, below). Their response to this question serves as the decision made by the human-AI team. In each phase, we control the distribution of gender and AI predictions. The participant sees applications from 2 women who are predicted as “Complete” and 3 women who are predicted as “Late” and vice versa for men. (This is true in the underlying data even if the participant and the model do not directly see each applicant’s gender.) We hold this ratio constant to avoid any effect due to the gender distribution or the rejection rate observed by different participants. To discourage participants from making decisions without any consideration of the prediction and (when applicable) explanation, we ask participants for a free-text justification of why they agreed or disagreed with the model prediction (or was neutral) after they have chosen their prediction on selected profiles. We randomly select one application in each gender + prediction combination for collecting these free-text justifications. These justifications also help us qualitatively assess the reasoning behind participants’ decisions. Further, to help filter out low-quality responses, participants are shown an attention check question, asking them to recall the previous AI prediction (Figure 9 in the Appendix) after seeing the first applicant in P1.
@@ 4.2. Procedure
Before proceeding to P2, participants may be shown general explanatory materials or specific disclosures on model bias and feature correlations. In the model bias disclosure (2(a)), participants are told that the model they saw in P1 had a low demographic parity (below $80\%$) (see §5.1 for details about demographic parity). In the correlation disclosure 2(b), participants are told the correlation between each university and gender in the model’s training data. The bias disclosure is shown across conditions, whereas correlation disclosure is only shown in the proxy conditions with correlation disclosure. In the proxy conditions without correlation disclosure, participants are only told that models can rely on proxy features to make biased predictions, without specifying the correlation between gender and university. This is done to make participants aware of potential biases without explicitly disclosing the correlations. Based on the disclosures seen, participants are asked up to two comprehension questions (Figure 13 in the Appendix). All are asked whether the model’s demographic parity was above $80\%$. Those who received correlation disclosure are asked to select one university that is highly associated with women. Note that before or during the first phase of the task, participants are never encouraged or primed to think about fairness explicitly. We only refer to fairness directly after phase 1. This allows us to measure how well participants can notice, or account for, unfairness when they aren’t explicitly told to look out for it in phase 1. Following this, in phase 2, we can then measure how participants perceive and account for unfairness when they know it is a salient concern.
@@
The three surveys (S0, S1, S2) aim to capture participants’ trust and fairness perceptions. All surveys include questions asking participants to rate their level of agreement with statements relating to trust (on a scale of $1$-$5$)  (Hoffman et al., 2018). In S0, participants are asked about their trust in AI systems generally, assessing their dispositional trust (Figure 14 in the Appendix). In S1 and S2, participants are asked about their trust in the system presented in the task phases, assessing their learned trust in the AI system that they interact with in the study (Figure 15 in the Appendix). In the post-task surveys (S1 and S2), alongside trust-related questions, participants are also asked about their perception of the fairness of the system they have been interacting with (whether “the AI system was fair across different genders”). Additionally, participants are asked the reason(s) that led to the their disagreements with AI such as the explanations including irrelevant features or the decisions being unfair towards applicants of different genders.
@@
In P0, participants are acclimatized to the task with a full tutorial example. They are shown one tutorial example with a walk-through of the task, the AI decisions and explanations (when applicable). Then they are shown warm-up examples. In the conditions without explanation, they are shown two example with no AI prediction or explanation. In the conditions with explanations, they are first shown a version of this example with no AI prediction or explanation. This is designed to encourage participants to properly engage with the features present. Second, they are shown the same example with the AI feature explanation (still without any prediction) as this setting has been shown to benefit decision quality and support learning by encouraging participants to cognitively engage with explanations (Gajos and Mamykina, 2022).
@@ 5. System Overview
We conduct our study using model predictions and explanations from logistic regression models trained on partially synthetic micro-lending data. Since the participant’s perceptions of how the model is interacting with the profile features is key to answering our research questions, we want to avoid any potential confounding effects from using artificial or Wizard-of-Oz model explanations, or entirely synthetic data. The scenario of predicting whether an applicant will complete microloan repayment on time or will be late is one that our participants will likely be sufficiently familiar with to have reasonable prior intuitions about what features are relevant. A challenge is that under US law, protected features like gender cannot be considered when making loan allocation decisions (Reserve, 2006) and therefore is not in the dataset that we consider. For this reason, we augment our data with a synthetic “gender” feature which we correlate with outcome to induce model bias. We also generate a proxy feature, university, which allows us to finely control the level of correlation between the proxy and gender.
@@
Our loan prediction data comes from a modified set of microloans from the website Prosper.1 The original dataset contains $79$ features of microloans including their status (completed, past due, etc). We group the loan statuses into “Complete” (including “Final Payment in Progress”), “Late” (including “Defaulted” and “Charged-Off”), or “Other” (including “Current” or “Canceld”). We keep the $\sim$14000140001400014000 profiles with “Complete” or “Late” statuses (with a $7$:$3$ train-test split). This grouped loan status is the feature that the participants and the model will predict. As showing all $79$ features to the participant may be overwhelming, (Poursabzi-Sangdeh et al., 2021) we select $5$ features (the original amount of the loan, the category of the listing, the applicant’s occupation and employment status, and their state of residence) that are both important to loan prediction and are likely interpretable by a layperson. As described above, we generate values for our protected characteristic (binary gender) synthetically. The existing applicants are assigned a gender such that women “Complete” vs are “Late” in repayment with a $2$:$3$ ratio and vice versa for men. This simulates historically biased data which will cause the model to associate femaleness with being late on loans and maleness with completing them. Using the generated “gender” feature, we further generate the proxy feature (university). We include co-ed and women’s colleges, setting the joint distribution of gender and university such that most co-ed universities have relatively balanced gender ratios (See 2(b)). For women’s colleges, the distributions reflect real-life statistics. We choose exclusively liberal arts colleges with similar US News rankings [^6] https://www.usnews.com/best-colleges/rankings/national-liberal-arts-colleges to avoid confounding due to the effect of perceptions of liberal arts vs non-liberal arts schools and perceptions of school rankings. Since, in our biased dataset, gender is correlated with outcome and, of course, the existing features are correlated with outcome, all features may be weakly correlated with gender. To confirm that university is the only strong proxy in our data, we compare the correlation of each categorical and continuous feature with gender. For continuous features (and one-hot features of each university), we use Pearson’s r coefficient. We find that the women’s colleges have at least an absolute correlation of $0.273$ across Proxy conditions, whereas the maximum absolute correlation for other continuous features is $0.014$, which is much lower. Similarly, for categorical features, we use Cramer’s V, finding that the university feature has at least an absolute correlation of $0.417$ while the maximum absolute correlation for the remaining categorical features is $0.082$, which is also lower. Overall, we see that university (especially women’s colleges) has a much stronger correlation with gender than any other feature shown to the participants.

### v2

# The Impact of Explanations on Fairness in Human-AI Decision-Making: Protected vs Proxy Features

@@ Abstract.
AI systems have been known to amplify biases in real-world data. Explanations may help human-AI teams address these biases for fairer decision-making. Typically, explanations focus on salient input features. If a model is biased against some protected group, explanations may include features that demonstrate this bias, but when biases are realized through proxy features, the relationship between this proxy feature and the protected one may be less clear to a human. In this work, we study the effect of the presence of protected and proxy features on participants’ perception of model fairness and their ability to improve demographic parity over an AI alone. Further, we examine how different treatments—explanations, model bias disclosure and proxy correlation disclosure—affect fairness perception and parity. We find that explanations help people detect direct but not indirect biases. Additionally, regardless of bias type, explanations tend to increase agreement with model biases. Disclosures can help mitigate this effect for indirect biases, improving both unfairness recognition and decision-making fairness. We hope that our findings can help guide further research into advancing explanations in support of fair human-AI decision-making.
@@
journalyear: 2024copyright: rightsretainedconference: 29th International Conference on Intelligent User Interfaces; March 18–21, 2024; Greenville, SC, USAbooktitle: 29th International Conference on Intelligent User Interfaces (IUI ’24), March 18–21, 2024, Greenville, SC, USAdoi: 10.1145/3640543.3645210isbn: 979-8-4007-0508-3/24/03ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Machine learning
@@ 1. Introduction
Improving the fairness and trustworthiness of AI systems is often cited as a goal of explainable AI (XAI) (e.g., Doshi-Velez and Kim, 2017; Lipton, 2018; Barredo Arrieta et al., 2020; Das and Rad, 2020; Langer et al., 2021; Du et al., 2021; Warner and Sloan, 2021). Research in XAI aims to improve fairness in human-AI decision-making by providing insights into model predictions, and thereby allowing humans to understand and correct for model biases. On the other hand, in the context of human-AI decision-making, previous work has noted that humans often over-rely on AI predictions, and explanations can exacerbate this concern (Buçinca et al., 2021). This is especially troubling if the underlying model contains systematic biases, which may go unnoticed even when teamed with a human. For the human-AI team to succeed, the human needs to be able to determine when to rely on or override potentially biased AI predictions. Previous work has shown that explanations can help human-AI teams alleviate model biases when those biases depend directly on protected attributes (Dodge et al., 2019; Schoeffer et al., 2022), but little is known in the very common case that protected attributes are not explicitly included, and rather the features used for prediction contain proxies thereof (e.g., zip code for race, length of credit for age, and university attended for gender). In particular, it may be difficult for humans to identify and resolve biased model predictions based on the proxy features present in real-world data, even when explanations are provided.
In this work, we study whether explanations can help people to identify model biases and to calibrate their reliance on a biased AI model. We extend work in this space by moving beyond direct biases that are revealed through the use of protected (i.e., sensitive) to indirect biases that are revealed through proxy features that may be less obvious to a human. Further, we examine whether explicitly disclosing model biases and correlations between the proxy and protected features can help humans calibrate their trust in a biased model. Our study aims to evaluate whether explanations can directly help notice model biases, even when the biases are obfuscated by the presence of proxy features and whether explanations can help users correct model biases when they are known to be present, through the use of bias disclosure and correlation disclosure. We study the effect of these treatments (explanations, model bias disclosure, and proxy correlation disclosure) on the fairness, including fairness perception and fairness in decision-making (measured by group-wise parity), as well as the accuracy of the decisions made by human-AI teams.
@@ 2. Background and Related Work
Both models and humans can be biased. Humans are known to exhibit many implicit and unconscious biases (Greenwald and Banaji, 1995). For instance,  Bertrand and Mullainathan (2004) find that an applicant with a “White-sounding name” on a resume that is otherwise identical to a resume with an “African-American-sounding name” is more likely to receive an interview callback.
@@
This paper explores how humans interact with predictions from a biased model, wherein AI systems may be able to uncover helpful patterns in existing data and humans may be able to apply their contextual understanding and societal awareness to contribute to correcting these biases within the model.
@@
Building upon previous research, this paper investigates the impact of explanations on the behavior of a human-AI team, especially their influence on the fairness of human-AI decisions in cases where the underlying model exhibits bias. In addition to explanations, we draw inspiration from work considering other methods of improving transparency in human-AI decision-making (e.g., Rechkemmer and Yin, 2022; Dzindolet et al., 2003; Chiang and Yin, 2021), exploring the implications of disclosing model bias and the correlation between protected and proxy features on the overall fairness of a human-AI team.
@@
As self-reported perceptions do not always align with observed behaviors in human-AI decision-making (Buçinca et al., 2020; Papenmeier et al., 2022), recent work has begun to expand out of fairness perceptions and into observed fairness in decision-making (Schoeffer et al., 2022; Wang et al., 2023). Schoeffer et al. (2022) study how explanations can help users appropriately rely on potentially unfair AI predictions. They find that explanations that highlight protected features negatively affect fairness perceptions and that decreases in fairness perception are associated with an increase in overrides of AI predictions, even on examples where this override is detrimental to the fairness of the human-AI team. Wang et al. (2023) study the effects of the level of model bias and the presence of explanation on the fairness of human decisions. They find that explanations lead participants to make more unfair decisions, even when participants were no longer given access to model predictions or explanations.
@@ 3. Research Questions
We study the effect of explanations and disclosures in improving the fairness perception and fairness of decisions made by human-AI teams. In our study, model biases can be direct: stemming from the protected feature (gender), or indirect: stemming from a proxy feature (university) that is correlated with the protected feature. For explanation, we consider an input-influence explanation of how each feature contributed to the AI’s prediction. For disclosures, participants may be told about the demographic parity (described in §5.1) of the system (model bias disclosure) and the strength of correlation between the proxy and protected features (proxy correlation disclosure; see §4.1). Our study addresses the following research questions:
@@
RQ5: Does dispositional trust (“an individual’s enduring tendency to trust automation” (Hoff and Bashir, 2015; Marsh and Dibben, 2003)) affect decision-making and fairness perceptions when working with models exhibiting direct or indirect bias?
RQ6: Do explanations and disclosures affect self-reported learned trust (based on “past experience or the current interaction” (Hoff and Bashir, 2015; Marsh and Dibben, 2003)) in models exhibiting direct or indirect bias?
In our study, we vary conditions based on the directness of bias, whether explanations are shown, and the kind of disclosure the participant receives. We consider six conditions. In the first three conditions, we do not show participants the explanations. Here, we consider one Protected condition with bias disclosure, and two Proxy conditions: one Proxy with correlation disclosure and one Proxy without correlation disclosure. We similarly consider three conditions with explanations allocating the biased feature and disclosure types in the same fashion.
We assess the effect of explanations (RQ1a) by comparing conditions with and without explanations (before any disclosures) in a between-subjects analysis. We assess the effect of disclosures without explanations (RQ2a) and the effect of disclosures with explanations (RQ3a) in a within-subject analysis comparing fairness perceptions and human-AI decision pre- and post-disclosures. This allows us to study how disclosures may help participants identify model biases over what is apparent before any disclosures. These first three effects are summarized in Figure 1. Lastly, we assess the effect of explanations and disclosures jointly (RQ4a) by comparing conditions in which participants are not shown explanations pre-disclosures with conditions in which participants are shown explanations post-disclosures. These experiments are repeated for protected and proxy conditions to assess the differences in interventions therein (RQ1-4b).
Within SubjectsBetween Subjects Phase 1 With Explanations Phase 2 With Explanations Phase 1 Without Explanations Phase 2 Without Explanations Effect of Explanations Alone (§6.1) Effect of Disclosure without Explanations (§6.2) Effect of Disclosure with Explanations (§6.3)
@@ 4. Study Design
To answer the research questions posed in §3, we study decisions made by human-AI teams. [^2] This study design is IRB approved. In this study, the AI teammate is a classification model trained on partially-synthetic data in the context of loan prediction. We choose the task of loan prediction from a micro-lending platform as it is a decision-making task performed by laypeople which means that crowd-workers are more likely to have intuitions about the task and the features used in predictions. In our study, participants are shown either the protected feature of binary gender [^3] We only consider binary gender in this study. Since each participant sees only a handful of examples per task phase, it would be difficult to both show the participants a statistically realistic number of non-binary applicants and get a good sense of how participants handle anti-trans model bias. We leave this for future work. or the proxy feature of university.
4.1. Procedure
Our study procedure consists of three surveys (S0, S1, S2), one tutorial and warm-up phase (P0), two task phases (P1, P2), and a disclosure interlude (D) ordered as shown in Figure 2.
S0 P0 P1 S1 D P2 S2 Initial survey: Dispositional trust Instructions and warm-up Refer to caption (×$\times 10$) Survey: Learned trust and fairness perceptions Disclosures: model bias and (according to condition) feature correlation Refer to caption (×$\times 10$) Survey: Learned trust and fairness perceptions
Figure 2. Order of study phases.
Task Phases
In each task phase (P1, P2), the participant is shown 10 profiles of loan applicants: their features and the overall AI prediction. Depending on the condition, they may or may not be shown an explanation of the AI prediction (Figure 3 left and right, respectively). This profile will, according to the condition, include either a “gender” or a “university” feature but not both. Participants are asked to mark on a five-point-scale whether they think the applicant will complete their loan on time or be late in repaying their loan (Figure 3, below). Their response to this question serves as the decision made by the human-AI team.
In each phase, we control the distribution of gender and AI predictions. The participant sees applications from 2 women who are predicted as “Complete” and 3 women who are predicted as “Late” and vice versa for men. (This is true in the underlying data even if the participant and the model do not directly see each applicant’s gender.) We hold this ratio constant to avoid any effect due to the gender distribution or the rejection rate observed by different participants.
To discourage participants from making decisions without any consideration of the prediction and (when applicable) explanation, we ask participants for a free-text justification of why they agreed or disagreed with the model prediction (or was neutral) after they have chosen their prediction on selected profiles. We randomly select one application in each gender + prediction combination for collecting these free-text justifications. These justifications also help us qualitatively assess the reasoning behind participants’ decisions. Further, to help filter out low-quality responses, participants are shown an attention check question, asking them to recall the previous AI prediction (Figure 9 in the Appendix) after seeing the first applicant in P1.
@@ 4.1. Procedure
Before proceeding to P2, participants may be shown general explanatory materials or specific disclosures on model bias and feature correlations. In the model bias disclosure (3(a)), participants are told that the model they saw in P1 had a low demographic parity (below $80\%$) (see §5.1 for details about demographic parity). In the correlation disclosure 3(b), participants are told the correlation between each university and gender in the model’s training data. The bias disclosure is shown across conditions, whereas correlation disclosure is only shown in the proxy conditions with correlation disclosure. In the proxy conditions without correlation disclosure, participants are only told that models can rely on proxy features to make biased predictions, without specifying the correlation between gender and university. This is done to make participants aware of potential biases without explicitly disclosing the correlations.
Based on the disclosures seen, participants are asked up to two comprehension questions (Figure 13 in the Appendix). All are asked whether the model’s demographic parity was above $80\%$. Those who received correlation disclosure are asked to select one university that is highly associated with women.
Note that participants are not encouraged or primed to consider fairness explicitly before or during the first phase of the task. We only refer to fairness directly after phase 1. This allows us to measure how well participants can notice, or account for, unfairness when they aren’t explicitly told to look out for it in phase 1. Subsequently, in phase 2, we can measure how participants perceive and account for unfairness when they know it is a salient concern.
@@
The three surveys (S0, S1, S2) aim to capture participants’ trust and fairness perceptions. All surveys include questions asking participants to rate their level of agreement with statements relating to trust (on a scale of $1$-$5$)  (Hoffman et al., 2018). In S0, participants are asked about their trust in AI systems generally, assessing their dispositional trust (Figure 14 in the Appendix). In S1 and S2, participants are asked about their trust in the system presented in the task phases, assessing their learned trust in the AI system that they interact with in the study (Figure 15 in the Appendix).
In the post-task surveys (S1 and S2), alongside trust-related questions, participants are also asked about their perception of the fairness of the system they have been interacting with (whether “the AI system was fair across different genders”). Additionally, participants are asked the reason(s) that led to their disagreements with AI such as the explanations including irrelevant features or the decisions being unfair towards applicants of different genders.
@@
In P0, participants are acclimatized to the task with a full tutorial example. They are shown one tutorial example with a walk-through of the task, the AI decisions and explanations (when applicable). Then, they are shown warm-up examples. In the conditions without explanation, they are shown two examples with no AI prediction or explanation. In the conditions with explanations, they are first shown a version of this example with no AI prediction or explanation. This is designed to encourage participants to properly engage with the features present. Second, they are shown the same example with the AI feature explanation (still without any prediction) as this setting has been shown to benefit decision quality and support learning by encouraging participants to cognitively engage with explanations (Gajos and Mamykina, 2022).
@@ 5. System Overview
We conduct our study using model predictions and explanations from logistic regression models trained on partially synthetic micro-lending data. [^6] https://github.com/ctbaumler/protected-vs-proxy Since the participant’s perceptions of how the model is interacting with the profile features is key to answering our research questions, we want to avoid any potential confounding effects from using artificial or Wizard-of-Oz model explanations, or entirely synthetic data.
The scenario of predicting whether an applicant will complete microloan repayment on time or will be late is one that our participants will likely be sufficiently familiar with to have reasonable prior intuitions about what features are relevant. A challenge is that under US law, protected features like gender cannot be considered when making loan allocation decisions (Reserve, 2006) and, therefore, is not in the dataset that we consider. For this reason, we augment our data with a synthetic “gender” feature, which we correlate with outcome to induce model bias. We also generate a proxy feature, university, which allows us to finely control the level of correlation between the proxy and gender.
@@
Our loan prediction data comes from a modified set of microloans from the website Prosper.1 The original dataset contains $79$ features of microloans including their status (completed, past due, etc). We group the loan statuses into “Complete” (including “Final Payment in Progress”), “Late” (including “Defaulted” and “Charged-Off”), or “Other” (including “Current” or “Canceld”). We keep the $\sim$14000140001400014000 profiles with “Complete” or “Late” statuses (with a $7$:$3$ train-test split). This grouped loan status is the feature that the participants and the model will predict. As showing all $79$ features to the participant may be overwhelming, (Poursabzi-Sangdeh et al., 2021) we select $5$ features (the original amount of the loan, the category of the listing, the applicant’s occupation and employment status, and their state of residence) that are both important to loan prediction and are likely interpretable by a layperson.
As described above, we synthetically generate values for our protected characteristic (binary gender). The existing applicants are assigned a gender in such a way that the ratio for “Complete” vs “Late” outcomes is $2$:$3$ for female applicants and vice versa for male applicants. This simulates historically biased data, which will cause the model to associate femaleness with being late on loans and maleness with completing them.
Using the generated “gender” feature, we further generate the proxy feature (university). We include co-ed and women’s colleges, setting the joint distribution of gender and university such that most co-ed universities have relatively balanced gender ratios (See 3(b)). For women’s colleges, the distributions reflect real-life statistics. We choose exclusively liberal arts colleges with similar US News rankings [^7] https://www.usnews.com/best-colleges/rankings/national-liberal-arts-colleges to avoid confounding due to the effect of perceptions of liberal arts vs non-liberal arts schools and perceptions of school rankings.
Since, in our biased dataset, gender is correlated with outcome and, of course, the existing features are correlated with outcome, all features may be weakly correlated with gender. To confirm that university is the only strong proxy in our data, we compare the correlation of each categorical and continuous feature with gender. For continuous features (and one-hot features of each university), we use Pearson’s r coefficient. We find that the women’s colleges have at least an absolute correlation of $0.273$ across Proxy conditions, whereas the maximum absolute correlation for other continuous features is $0.014$, which is much lower. Similarly, for categorical features, we use Cramer’s V, finding that the university feature has at least an absolute correlation of $0.417$ while the maximum absolute correlation for the remaining categorical features is $0.082$, which is also lower. Overall, we see that university (especially women’s colleges) has a much stronger correlation with gender than any other feature shown to the participants.